Circular No. 6028

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) T. Rettig, J. Hahn, S. Tegler, M. Mumma and M. DiSanti (cf. IAUC

6019) report a “last look” detection of fragment K = 12 out at a distance of ∼ 20 jovian radii in a Hubble Space Telescope WFPC image on July 18.554 UT, 21 hours prior to impact. Comparison with a June 26 image shows that the elongation of the inner coma has become even more pronounced. This elongation is parallel to the jovian direction and also appears slightly brighter or more extended toward Jupiter. Preliminary analysis of the July 18 image shows only the region within ∼ 0′′.1 of the central pixel to have circular symmetry. At 0′′.3–0′′.4 from this pixel the surface brightness aspect ratio has increased to ∼ 2:1 and at 0′′.7–0′′.8 to 2.5–3:1.

The K impact has been documented in reports received via the SL9 message center. SPIREX (cf. IAUC 6026) imaged the impact over the limb at 2.36 μm on July 19.440 UT, the peak brightness appearing comparable to that of the G impact. At Okayama Astrophysical Observatory (cf. IAUC 6025) an initial 2.36-μm flash was detected on July 19.434, followed by an increase on July 19.438 that reached a maximum around 19.443 of perhaps > 20 times the brightness of Io. M. A’Hearn, University of Maryland, reports that R filter observations at the Perth Observatory during July 19.417–19.437 showed no hint of a flash reflected from Europa, then in eclipse; beginning July 19.441 methane 893-nm filter observations showed a definite prominence above the apparent limb of Jupiter having approximately the same surface brightness as other portions of that latitude belt. CASPIR (cf. IAUC 6024) confirmed that that there was no reflected emission from Europa and noted the first indication of a plume at 2.34 μm on July 19.432; on 19.434 a bright fireball was detected and seen to grow to twice the size of the nearest remnant impact site by 19.435, after which it remained constant until 19.439, when an intense central core brightened enough to cast diffraction spikes across the array; as the site rotated into view the fireball increased in brightness, and by July 19.440 it was estimated as slightly brighter than the peak brightness of impact G yesterday.

A message-center report by E. Lellouch, M. Festou and G. Paubert mentions that IRAM 30-m observations of site G on July 18.760 and 18.861 UT (∼ 25 min integration in each case) shows the CO 230-GHz line, ∼ 0.6 K contrast, 2.5 km/s FWHM wide and integrated intensity ∼ 1.5 K km/s. A marginal detection at site H around July 18.819 possibly showed the line with an intensity of ∼ 0.3 K km/s. Further high-resolution (100 kHz or so) monitoring is urged.

1994 July 19 Brian G. Marsden